OpenStudy (anonymous):

Help Me please :c Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem. Can you give me the answer and how you got it for me, so I can see if I get the same thing, thank you :3

OpenStudy (anonymous):

What does mass media mean?

OpenStudy (ikram002p):

assume x is flowers num profit=x(1 flower price - 1 flower cost) 500=x((0.5+0.5×40%)-0.5) 500=x(0.2) x=2500 flowers

OpenStudy (anonymous):

I got the same thing, thank you so much :D

OpenStudy (anonymous):

But, how do you put that into a sentence? Can you help me? @ikram002p

OpenStudy (ikram002p):

here is the sentence assume x is flowers num profit=x(1 flower price - 1 flower cost)

OpenStudy (anonymous):

…Ok, I can make a good sentence with that, thank you some much for the help :3

OpenStudy (ikram002p):

wlc :)